1. Jennie dropped a stone from the top of the blackpink house which is 400 m high.


  1. How long will it take for the stone to reach the ground?
1
Expert's answer
2021-09-08T16:13:03-0400

Given:

"H=400\:\rm m"


The time of free falling of the stone

"t=\sqrt{2H/g}=\sqrt{2*400\:\rm m/9.8\:\rm m/s^2}=9.0\:\rm s"